# Search for Past Stellar Encounters and the Origin of 3I/ATLAS

### v1

@@ 3.1 Preliminary Selection of Encounters
We first select encounter candidates with nominal distance $d^{\mathrm{lin}}_{\mathrm{enc}}<50$ pc, yielding $\sim$350,000 candidates. We perform orbital integration using the galpy package (Bovy, 2015), over a $\pm 100$ Myr time span with a timestep of 2000 yr. After computing discrete orbital trajectories, we apply cubic spline interpolation to extract the periastron parameters. We test orbital propagation with time steps of 2000, 1000, and 100 yr and find negligible differences in encounter results.
@@ 5 Cumulative Effects of Encounter Scattering
We analyze each encounter’s gravitational slingshot to evaluate whether 3I/ATLAS originated near the Sun’s location, namely about 8 kpc from the Galactic center and embedded in the thin disk, and subsequently was accelerated to its present high speed. We model 3I/ATLAS encounters by assuming that stars are distributed in a cylindrical tube of maximum radius $d_{\mathrm{max}}$ around the object’s trajectory, and the encounter rate is uniformly distributed over time $T$. Following (Feng & Bailer-Jones, 2014), the unnormalized probability density is:
@@ 6 Discussion and Conclusion
To clarify the magnitude of encounter perturbations to 3I/ATLAS, we use an analytic impulse model to calculate the deflection angle of G 137-55. Even though G 137-55 yields the largest significance of perturbation $g$, the resulting deflection angle is only a few arcseconds. The current precision of star astrometry and radial velocity contributes much more uncertainty than scattering with the stars, especially when the encounter time $|t^{\mathrm{}}_{\mathrm{enc}}|$ is large. On Gyr timescales, the cumulative encounter-driven changes in both speed and direction are found to be very weak for 3I/ATLAS’s current speed. Using empirical fits to the velocity dispersions of thin-disk and thick-disk stellar populations, the velocity of 3I/ATLAS is far more probable under the thin-disk distribution. Moreover, because thin-disk stars dominate the local number density serving as a natural prior, the posterior probability that 3I/ATLAS is ejected from the thin disk is further enhanced.
Additionally, when candidate encounters are derived from the Gaia DR3, many of the stars lack robust spectral types, age estimates, or reliable masses. This scarcity of stellar parameters further hinders the identification of host stars for 3I/ATLAS at the current stage. As additional interstellar objects are discovered, a comprehensive evaluation system for determining host stars of interstellar objects will be necessary for deepening our understanding of planetary systems beyond the Solar System.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China (Nos. 2024YFA1611801 and 2024YFC2207800) and by the National Natural Science Foundation of China (NSFC, Grant No. 12473066). It was also supported by the Shanghai Jiao Tong University 2030 Initiative and the China–Chile Joint Research Fund (CCJRF No. 2205). CCJRF is provided by the Chinese Academy of Sciences South America Center for Astronomy (CASSACA) and was established by the National Astronomical Observatories, Chinese Academy of Sciences (NAOC), and the Chilean Astronomy Society (SOCHIAS) to support China–Chile collaborations in astronomy. Additional support was provided by the Fundamental Research Funds for the Central Universities, the ”111 Project” (No. B20019), and the Shanghai Natural Science Foundation (Grant No. 19ZR1466800). Luyao was supported by the China Scholarship Council (CSC) in collaboration with the University of Leicester. This work is based on data from the European Space Agency (ESA) mission Gaia (www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; www.cosmos.esa.int/web/gaia/dpac/consortium).

### v2

@@ 3.1 Preliminary Selection of Encounters
We first select encounter candidates with nominal distance $d^{\mathrm{lin}}_{\mathrm{enc}}<50$ pc, yielding $\sim$350,000 candidates. As shown by Bailer-Jones (2015b), adopting a sufficiently large threshold on linear encounter distance $d^{\mathrm{lin}}_{\mathrm{enc}}$ does not miss targets that would qualify as genuine close encounters in subsequent nonlinear propagation.
We perform orbital integration using the galpy package (Bovy, 2015), over a $\pm 100$ Myr time span with a timestep of 2000 yr. After computing discrete orbital trajectories, we apply cubic spline interpolation to extract the periastron parameters. We test orbital propagation with time steps of 2000, 1000, and 100 yr and find negligible differences in encounter results.
@@ 5 Cumulative Effects of Encounter Scattering
We analyze each encounter’s gravitational slingshot to evaluate whether 3I/ATLAS originated near the Sun’s location, namely about 8 kpc from the Galactic center and embedded in the thin disk, and subsequently was accelerated to its present high speed. We model 3I/ATLAS encounters by assuming that stars are distributed in a cylindrical tube of maximum radius $d_{\mathrm{max}}$ around the object’s trajectory, and the encounter rate is uniformly distributed over time $T$. Following Feng & Bailer-Jones (2014), the unnormalized probability density is:
@@ 6 Discussion and Conclusion
To clarify the magnitude of encounter perturbations to 3I/ATLAS, we use an analytic impulse model to calculate the deflection angle of G 137-55. Even though G 137-55 yields the largest significance of perturbation $g$, the resulting deflection angle is only a few arcseconds. The current precision of star astrometry and radial velocity contributes much more uncertainty than scattering with the stars, especially when the encounter time $|t^{\mathrm{}}_{\mathrm{enc}}|$ is large. On Gyr timescales, the cumulative encounter-driven changes in both speed and direction are found to be very weak for 3I/ATLAS’s current speed. These considerations likewise support using nonlinear orbit integrations in a smooth galactic potential as a sufficiently accurate alternative to, in principle, full $N$-body modeling with stars and the galactic potential. At the million-body level, such $N$-body calculations are computationally prohibitive for our used time steps. On a 100 Myr timescale, impulses from close encounters and second-order star–star terms produce only negligible perturbations to the trajectory of 3I/ATLAS.
On a 100 Myr timescale, we adopt a static galactic potential. The Sun’s orbital period is around 220 Myr, and key time-dependent drivers evolve on longer or comparable scales. For example the vertical phase-mixing timescale associated with the Gaia phase spiral is approximately 0.3 to 0.9 Gyr (Grand et al., 2012; Hunt et al., 2018; Reid et al., 2019). Therefore, treating the potential as approximately time-independent on this timescale is a common approximation. Using empirical fits to the velocity dispersions of thin-disk and thick-disk stellar populations, the velocity of 3I/ATLAS is far more probable under the thin-disk distribution. Moreover, because thin-disk stars dominate the local number density serving as a natural prior, the posterior probability that 3I/ATLAS is ejected from the thin disk is further enhanced.
Additionally, when candidate encounters are derived from the Gaia DR3, many of the stars lack robust spectral types, age estimates, or reliable masses. This scarcity of stellar parameters further hinders the identification of host stars for 3I/ATLAS at the current stage. Machine-learning methods can be used to integrate the currently observed properties of ISOs to better infer their host stars and birth environments. In particular, by simulating both the ejection of ISOs and the physical processing they undergo during interstellar travel like cosmic-ray irradiation and space weathering, observable features such as kinematics, colors, and elemental abundances enable models to learn nonlinear correlations and to represent full probability distributions in a high-dimensional feature space. Compared with traditional orbit-integration pipelines, these models are designed to improve origin inference for ISOs. As additional interstellar objects are discovered, a comprehensive evaluation system for determining host stars of interstellar objects will be necessary for deepening our understanding of planetary systems beyond the Solar System.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China (Nos. 2024YFA1611801 and 2024YFC2207800) and by the National Natural Science Foundation of China (NSFC, Grant No. 12473066). It was also supported by the Shanghai Jiao Tong University 2030 Initiative and the China–Chile Joint Research Fund (CCJRF No. 2205). CCJRF is provided by the Chinese Academy of Sciences South America Center for Astronomy (CASSACA) and was established by the National Astronomical Observatories, Chinese Academy of Sciences (NAOC), and the Chilean Astronomy Society (SOCHIAS) to support China–Chile collaborations in astronomy. Additional support was provided by the Fundamental Research Funds for the Central Universities, the ”111 Project” (No. B20019), and the Shanghai Natural Science Foundation (Grant No. 19ZR1466800). This project was also supported by the Office of Science and Technology, Shanghai Municipal Government (grants 24DX1400100 and ZJ2023-ZD-001), and by the SJTU-Warwick Joint Seed Fund 2024/25-Round 5). Luyao was supported by the China Scholarship Council (CSC) in collaboration with the University of Leicester. This work is based on data from the European Space Agency (ESA) mission Gaia (www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; www.cosmos.esa.int/web/gaia/dpac/consortium).